Trying to figure out this story problem and I am having some difficulty.
A disc jockey must play 16 commercial spots during 1 hour of a radio show. Each commercial is either 30 seconds or 60 seconds long. If the total commercial time during 1 hour is 13 mins. How many 30-second commercials were played that hour? And how many 60-second commercials?
Thanks :o)

Answer by mananth(16946)

let 30 second commercial be x
60 second commercial will be 16-x
13 minutes = 13*60 seconds = 780
..
30x+60(16-x)= 780
30x+960-60x=780
-30x+960=780
add -960 to both sides
-30x+960-960=780-960
-30x=-180
divide by -30
x=-180/-30
x= 6 which is 30 second commercial
16-x = 16-6= 10 60 second commercials
..
CHECK
6*30+10*60=780 seconds= 13 minutes